# Investigating symptom duration using current status data: a case study of post-acute COVID-19 syndrome

### v1

@@ Abstract
For infectious diseases, characterizing symptom duration is of clinical and public health importance. Symptom duration may be assessed by surveying infected individuals and querying symptom status at the time of survey response. For example, in a SARS-CoV-2 testing program at the University of Washington, participants were surveyed at least 28 days after testing positive and asked to report current symptom status. This study design yielded current status data: Outcome measurements for each respondent consisted only of the time of survey response and a binary indicator of whether symptoms had resolved by that time. Such study design benefits from limited risk of recall bias, but analyzing the resulting data necessitates specialized statistical tools. Here, we review methods for current status data and describe a novel application of modern nonparametric techniques to this setting. The proposed approach is valid under weaker assumptions compared to existing methods, allows use of flexible machine learning tools, and handles potential survey nonresponse. From the university study, we estimate that 19% of participants experienced ongoing symptoms 30 days after testing positive, decreasing to 7% at 90 days. Female sex, history of seasonal allergies, fatigue during acute infection, and higher viral load were associated with slower symptom resolution.
@@ Introduction
The exact time from infection until symptom resolution was not observed for any participant in the HCT long COVID study. Instead, for each survey respondent, in addition to baseline covariate information collected at or before the test date, the available data consist of the time elapsed from positive SARS-CoV-2 test until survey response — referred to as the survey response time — and a binary variable indicating whether the respondent’s symptoms had resolved as of the survey response time — referred to as the event indicator. Each participant was observed only at a single survey response time. Response times in this study ranged from 28 days (the earliest time after a positive test at which surveys were sent) to 286 days. Because by design investigators recorded only the current status of a participant’s symptom persistence at their response time, this sampling scheme is referred to as current status sampling. Current status data, which represent an extreme form of interval censoring, arise in a number of epidemiological settings, including partner studies of HIV infection and cross-sectional studies for estimating the age of onset of non-fatal diseases, and require the use of specialized statistical techniques [8]. Study designs analogous to that of the HCT long COVID survey may be attractive because they are simple to implement, economical, and most importantly, less susceptible to recall bias than retrospective designs in which participants are asked to recall symptom status at a specific date in the past [9]. As with other survey designs, nonresponse is a potential concern: When the study concluded in June 2023, a large fraction of participants had not yet responded to the long COVID survey.
@@
Interest lies in (1) estimating the distribution of the time-to-symptom-resolution variable and (2) identifying baseline patient characteristics associated with time to resolution, that is, risk factors for delayed symptom resolution. To achieve (1), we show here that recent methodological advances we have made in causal inference under shape constraints can be adapted to make statistical inference on the time-to-symptom-resolution distribution. In particular, this approach allows us to use flexible machine learning tools and thereby minimize the need for restrictive modeling assumptions. To our knowledge, the link between the methodology we build upon here, termed causal isotonic regression (CIR), and the analysis of current status data has not been made before; we aim to highlight the utility of this approach. To achieve (2), we make use of existing regression methods for current status data, which we briefly describe in the Methods section below, with further details given in the Supplementary Material. Our chosen approaches to both (1) and (2) rely on identical assumptions regarding the survey response process, namely, that survey response times are uninformative for symptom resolution within strata of measured baseline covariates. In addition to illustrating statistical tools tailored to the analysis of current status data, we provide substantive results regarding persistent COVID-19 symptoms in the HCT population during the early Omicron period of the SARS-CoV-2 pandemic.
@@ Methods
We begin by describing in general the current status data structure we are considering, using the HCT study for concreteness. For each recipient of the long COVID survey, a collection $W$ of baseline covariates, observed at or before the time of a positive SARS-CoV-2 test, were recorded. Estimating the distribution of symptom duration $T$, the time elapsed between infection and symptom resolution, is challenging because (1) $T$ is never observed directly and (2) the study is terminated before all participants have responded to the survey.
In an ideal study with unlimited follow-up, we would observe for each participant the survey response time $Y_{*}$ — the time elapsed between positive SARS-CoV-2 test and survey response — and the event indicator $\Delta_{*}$ — a participant’s symptom status at the time of survey response — with $\Delta_{*}=1$ representing symptom resolution by time $Y_{*}$ and $\Delta_{*}=0$ otherwise. We denote by $c_{0}$ the investigator-determined time (since the positive test) at which follow-up ends for a participant, and consider as a nonrespondent any participant without a response by time $c_{0}$. We set $Y=\min(Y_{*},c_{0})$ as the smaller of the survey response time and $c_{0}$, and denote by $\Delta=\mathbbm{1}(Y_{*}<c_{0})\Delta_{*}$ the observed response indicator, which equals 1 if a participant responded and reported resolution and 0 otherwise. The available data are considered to be $n$ independent and identically distributed observations $(W_{1},Y_{1},\Delta_{1}),(W_{2},Y_{2},\Delta_{2}),\ldots,(W_{n},Y_{n},\Delta_{n})$.
The two primary goals of our analysis, as described earlier, can be stated symbolically as using the available data to:
  1. 1.
    estimate the survival function of $T$;
  2. 2.
    estimate the association between baseline covariates $W$ and symptom duration $T$.
Below, we describe how these two goals may be achieved.
@@
For a given time $t$, we denote by $\psi(t)=P(T>t)$ the survival function of $T$ evaluated at $t$. This quantity represents the proportion of individuals testing positive for SARS-CoV-2 who experience symptom persistence beyond time $t$ days. Our aim is to estimate $\psi(t)$ over an interval $[t_{0},t_{1}]$ strictly contained in $(b_{0},c_{0})$, where $b_{0}$ = 28 days is the smallest observable response time. Estimating this function using current status data requires dedicated statistical tools, since failure to account for the sampling scheme can lead to misleading conclusions.
@@
Computation of the NPMLE involves performing an isotonic regression of $\Delta_{1},\Delta_{2},\ldots,\Delta_{n}$ on $Y_{1},Y_{2},\ldots,Y_{n}$ over the domain $[t_{0},t_{1}]$. Computation of the CIR estimator is more complex since it includes a step to disentangle the dependence between event and survey response times. This step involves estimation of two functions depending on baseline covariates and survey response times: (1) $\mu(y,w)=E(\Delta\,|\,Y=y,W=w)$, the conditional mean of $\Delta$ given $Y$ and $W$; and (2) $g(y,w)=\pi(y\,|\,w)/E\{\pi(y\,|\,W)\}$, a standardization of the conditional density $\pi$ of $Y$ given $W$. These functions play a role akin to those of the outcome regression and (standardized) propensity score in causal inference for continuous exposures. To implement the CIR procedure, estimates $\mu_{n}$ and $g_{n}$ of $\mu$ and $g$, respectively, are first constructed; this task can be accomplished with any off-the-shelf machine learning algorithm. Pseudo-outcomes $\Gamma_{n}(\Delta_{1},Y_{1},W_{1}),\Gamma_{n}(\Delta_{2},Y_{2},W_{2}),\ldots,\Gamma_{n}(\Delta_{n},Y_{n},W_{n})$, defined pointwise as
@@
We implemented the extended CIR method using flexible machine learning tools to estimate $\mu$ and $g$. To estimate $\mu$, we used the Super Learner [15, 16], a form of stacked regression, with the optimal linear combination of learners selected to minimize 10-fold cross-validated mean squared error — see Supplementary Material for algorithm library. To estimate $\pi$, we used the haldensify software package [17] with 10-fold cross-validation used to select optimal tuning parameters. The derivative of the survival curve, which is required for confidence interval construction, was estimated by numerical differentiation of a monotone Hermite spline approximation of the estimated survival curve. Confidence intervals with nominal 95% pointwise coverage were constructed using the quantiles of the Chernoff distribution.
@@ Results
After the removal of 284 individuals due to missing covariate information, the analysis cohort consisted of 3434 HCT study participants with positive SARS-CoV-2 tests prior to February 24, 2023. Of these, 1432 (41.7%) responded to the survey. Characteristics of the analysis cohort are summarized in Table 3. The median age was 23 years (IQR: 20–36), and approximately 64% of the participants reported female sex. Slightly over 33% of participants reported at least one baseline comorbidity, and roughly 52% were symptomatic at the time of testing. Most participants were not recently vaccinated. Survey respondents tended to be older than nonrespondents, with markedly higher prevalence of symptoms during the acute infection phase.
@@
Figure 1 shows the estimated proportion of participants with persistent COVID-19 symptoms at times from 30 to 90 days following the positive test, along with the distribution of survey response times. From the extended CIR method, estimated symptom persistence proportions at 30, 60, and 90 days after the positive test were 19.0% (95% CI: 13.6–25.0), 13.0% (9.4–16.7), and 6.9% (2.9–10.9), respectively.
Figure 1: Covariate-adjusted survival curves for time from positive test until symptom resolution, estimated using the proposed extended causal isotonic regression method (top panel), along with the distribution of survey response times between 28 and 95 days (bottom panel). The solid black line shows the estimated proportion with ongoing symptoms at times from 30 days to 90 days after the positive test. The black points show the estimated proportion with ongoing symptoms at 30, 60, and 90 days. The dashed lines represent a pointwise 95% confidence interval. The gray vertical lines denote 30, 60, and 90 days.
Table 4 contains the results of the Cox proportional hazards regression, with hazard ratios larger than 1 indicating faster symptom resolution. Participants identifying as male tended to experience faster symptom resolution, with an estimated hazard ratio for resolution of 1.21 (1.04–1.40, $P$ = 0.018). Conversely, participants reporting fatigue at baseline and participants with history of seasonal allergies tended to have longer lasting symptoms, with estimated hazard ratios of 0.77 (0.65–0.92, $P$ = 0.004) and 0.86 (0.74–1.00, $P$ = 0.050), respectively. In terms of viral load, participants with Ct values under 30 and participants who tested positive via rapid antigen tests had more persistent symptoms compared to participants with Ct values over 30, with estimated hazard ratios of 0.78 (0.64–0.96) and 0.75 (0.60–0.95), respectively ($P$ = 0.033).
@@ Discussion
In this article, we presented modern methodology for the analysis of current status data arising in epidemiological studies, with a particular focus on the novel extension and application of causal isotonic regression to this setting. When it is of interest to estimate the distribution of disease duration under current status sampling, CIR offers a flexible approach accommodating the use of off-the-shelf machine learning tools. Unlike the commonly used nonparametric maximum likelihood estimator, the validity of CIR relies only on conditional independence of survey response times and event indicators within strata of measured baseline covariates. In addition, we have shown how CIR can be implemented to properly account for survey nonresponse.

### v2

@@ Abstract
For infectious diseases, characterizing symptom duration is of clinical and public health importance. Symptom duration may be assessed by surveying infected individuals and querying symptom status at the time of survey response. For example, in a SARS-CoV-2 testing program at the University of Washington, participants were surveyed at least 28 days after testing positive and asked to report current symptom status. This study design yielded current status data: outcome measurements for each respondent consisted only of the time of survey response and a binary indicator of whether symptoms had resolved by that time. Such study design benefits from limited risk of recall bias, but analyzing the resulting data necessitates tailored statistical tools. Here, we review methods for current status data and describe a novel application of modern nonparametric techniques to this setting. The proposed approach is valid under weaker assumptions compared to existing methods, allows use of flexible machine learning tools, and handles potential survey nonresponse. From the university study, we estimate that 19% of participants experienced ongoing symptoms 30 days after testing positive, decreasing to 7% at 90 days. Female sex, history of seasonal allergies, fatigue during acute infection, and higher viral load were associated with slower symptom resolution.
@@ Introduction
The exact time from infection until symptom resolution was not observed for any participant in the HCT long COVID study. Instead, for each survey respondent, in addition to baseline covariate information collected at or before the test date, the available data consist of the time elapsed from positive SARS-CoV-2 test until survey response — referred to as the survey response time — and a binary variable indicating whether the respondent’s symptoms had resolved as of the survey response time — referred to as the event indicator. Each participant was observed only at a single survey response time. Response times in this study ranged from $28$ days (the earliest time after a positive test at which surveys were sent) to $286$ days. Because by design investigators recorded only the current status of a participant’s symptom persistence at their response time, this sampling scheme, which represents an extreme form of interval censoring, is referred to as current status sampling.
Current status data arise and have been analyzed in a number of epidemiological settings, including partner studies of HIV transmission [8], cross-sectional studies for estimating the age of onset of a non-fatal disease [9, 10], and demographic studies of age at weaning [11, 12, 13, 14, 15]. Designs analogous to the HCT long COVID study, in which participants are surveyed following a health-related event and queried regarding the presence of sequelae, have been used, for example, to study outcomes following radiation therapy [16], joint surgery [17, 18], diagnosis with light chain amyloidosis [19], onset of schizophrenia [20], and infection with Ebola virus [21]. In fact, previous studies of persistent COVID-19 symptoms have also used a similar design [22, 23, 24, 25]. Such designs may be appealing because they are simple to implement, economical, and most importantly, less susceptible to recall bias than retrospective designs in which participants are asked to recall symptom status at a specific date in the past [26]. Nevertheless, the presence of current status data may be easily missed; many of the aforementioned studies treated occurrence of the outcome as a binary variable, ignoring variation in response times. Recognizing the presence of current status data is vital, since their analysis requires specialized statistical techniques [27].
@@
Traditional analysis techniques for current status data require that survey response times be uninformative of the symptom resolution process. In this paper, we show that recent methodological advances we have made in causal inference under shape constraints can be adapted to make statistical inference on the time-to-symptom-resolution distribution allowing some response time informativeness. The proposed approach allows us to use flexible machine learning tools and thereby minimize the need for restrictive modeling assumptions. Furthermore, our method accounts for potential nonresponse, which is ubiquitous in survey studies such as the HCT long COVID study, where over half of survey recipients had not responded when the study concluded. To our knowledge, the link between the methodology we build upon here, termed causal isotonic regression (CIR), and the analysis of current status data has not been made before; we aim to highlight the utility of this approach. Using our method, we provide substantive results regarding persistent COVID-19 symptoms in the HCT population during the early Omicron period of the SARS-CoV-2 pandemic. We also investigate baseline patient characteristics associated with time to resolution, that is, risk factors for delayed symptom resolution, using existing regression methods for current status data.
@@ Methods
We begin by describing in general the current status data structure we are considering, using the HCT study for concreteness. For each recipient of the long COVID survey, a collection $W$ of baseline covariates, observed at or before the time of a positive SARS-CoV-2 test, were recorded. Estimating the distribution of symptom duration $T$, the time elapsed between infection and symptom resolution, is challenging because ($1$) $T$ is never observed directly and ($2$) the study is terminated before all participants have responded to the survey.
In an ideal study with unlimited follow-up, we would observe for each participant the survey response time $Y_{*}$ — the time elapsed between positive SARS-CoV-2 test and survey response — and the event indicator $\Delta_{*}$ — a participant’s symptom status at the time of survey response — with $\Delta_{*}=1$ representing symptom resolution by time $Y_{*}$ and $\Delta_{*}=0$ otherwise. We denote by $c_{0}$ the investigator-determined time (since the positive test) at which follow-up ends for a participant, and consider as a nonrespondent any participant without a response by time $c_{0}$. We set $Y=\min(Y_{*},c_{0})$ as the smaller of the survey response time and $c_{0}$, and denote by $\Delta=\mathbbm{1}(Y_{*}<c_{0})\Delta_{*}$ the observed response indicator, which equals $1$ if a participant responded and reported resolution and 00 otherwise. The available data are considered to be $n$ independent and identically distributed observations $(W_{1},Y_{1},\Delta_{1}),(W_{2},Y_{2},\Delta_{2}),\ldots,(W_{n},Y_{n},\Delta_{n})$.
In statistical terms, the primary goal of our analysis consists of using the available data to estimate the survival function of $T$. It may also be of interest to estimate the association between baseline covariates $W$ and symptom duration $T$. Below, we describe how these two goals may be achieved.
@@
For a given time $t$, we denote by $\psi(t)=P(T>t)$ the survival function of $T$ evaluated at $t$. This quantity represents the proportion of individuals testing positive for SARS-CoV-2 who experience symptom persistence beyond time $t$ days. Our aim is to estimate $\psi(t)$ over an interval $[t_{0},t_{1}]$ strictly contained in $(b_{0},c_{0})$, where $b_{0}=28$ days is the smallest observable response time. Estimating this function using current status data requires tailored statistical tools, since failure to account for the sampling scheme can lead to misleading conclusions.
@@
Computation of the NPMLE involves performing an isotonic regression of $\Delta_{1},\Delta_{2},\ldots,\Delta_{n}$ on $Y_{1},Y_{2},\ldots,Y_{n}$ over the domain $[t_{0},t_{1}]$. Computation of the CIR estimator is more complex since it includes a step to disentangle the dependence between event and survey response times. This step involves estimation of two functions depending on baseline covariates and survey response times: ($1$) $\mu(y,w)=E(\Delta\,|\,Y=y,W=w)$, the conditional mean of $\Delta$ given $Y$ and $W$; and ($2$) $g(y,w)=\pi(y\,|\,w)/E\{\pi(y\,|\,W)\}$, a standardization of the conditional density $\pi$ of $Y$ given $W$. These functions play a role akin to those of the outcome regression and (standardized) propensity score in causal inference for continuous exposures. To implement the CIR procedure, estimates $\mu_{n}$ and $g_{n}$ of $\mu$ and $g$, respectively, are first constructed; this task can be accomplished with any off-the-shelf machine learning algorithm. Pseudo-outcomes $\Gamma_{n}(\Delta_{1},Y_{1},W_{1}),\Gamma_{n}(\Delta_{2},Y_{2},W_{2}),\ldots,\Gamma_{n}(\Delta_{n},Y_{n},W_{n})$, defined pointwise as
@@
We implemented the extended CIR method using flexible machine learning tools to estimate $\mu$ and $g$. To estimate $\mu$, we used the Super Learner [33, 34], a form of stacked regression, with the optimal linear combination of learners selected to minimize $10$-fold cross-validated mean squared error — see Supplementary Material for the algorithm library. To estimate $\pi$, we used the haldensify software package [35] with $10$-fold cross-validation used to select optimal tuning parameters. The derivative of the survival curve, which is required for confidence interval construction, was estimated by numerical differentiation of a monotone Hermite spline approximation of the estimated survival curve. Confidence intervals with nominal $95$% pointwise coverage were constructed using the quantiles of the Chernoff distribution.
@@ Results
After the removal of $284$ individuals due to missing covariate information, the analysis cohort consisted of $3434$ HCT study participants with positive SARS-CoV-2 tests prior to February $24$, $2023$. Of these, $1432$ ($41.7$%) responded to the survey. Characteristics of the analysis cohort are summarized in Table 3. The median age was $23$ years (IQR: 2020202036363636), and approximately $64$% of the participants reported female sex. Slightly over $33$% of participants reported at least one baseline comorbidity, and roughly $52$% were symptomatic at the time of testing. Most participants were not recently vaccinated. Survey respondents tended to be older than nonrespondents, with markedly higher prevalence of symptoms during the acute infection phase.
@@
Figure 1 shows the estimated proportion of participants with persistent COVID-19 symptoms at times from 30 to 90 days following the positive test, along with the distribution of survey response times. From the extended CIR method, estimated symptom persistence proportions at $30$, $60$, and $90$ days after the positive test were $19.0$% (95% CI: 13.613.613.613.625.025.025.025.0), $13.0$% (9.49.49.49.416.716.716.716.7), and $6.9$% (2.92.92.92.910.910.910.910.9), respectively.
Figure 1: Standardized survival curve for time from positive test until symptom resolution, estimated using the proposed extended causal isotonic regression method (top panel), along with the distribution of survey response times between $28$ and $95$ days (bottom panel). The solid black line shows the estimated proportion with ongoing symptoms at times from $30$ days to $90$ days after the positive test. The black points show the estimated proportion with ongoing symptoms at $30$, $60$, and $90$ days. The dashed lines represent a pointwise 95% confidence interval. The gray vertical lines denote $30$, $60$, and $90$ days.
Table 4 contains the results of the Cox proportional hazards regression, with hazard ratios larger than $1$ indicating faster symptom resolution. Participants identifying as male tended to experience faster symptom resolution, with an estimated hazard ratio for resolution of $1.21$ (1.041.041.041.041.401.401.401.40, $P=0.018$). Conversely, participants reporting fatigue at baseline and participants with history of seasonal allergies tended to have longer lasting symptoms, with estimated hazard ratios of $0.77$ (0.650.650.650.650.920.920.920.92, $P=0.004$) and 0.86 (0.740.740.740.741.001.001.001.00, $P=0.050$), respectively. In terms of viral load, participants with Ct values under $30$ and participants who tested positive via rapid antigen tests had more persistent symptoms compared to participants with Ct values over $30$, with estimated hazard ratios of $0.78$ (0.640.640.640.640.960.960.960.96) and $0.75$ (0.600.600.600.600.950.950.950.95), respectively ($P=0.033$).
@@ Discussion
In this article, we presented modern methodology for the analysis of current status data arising in epidemiological studies, with a particular focus on the novel extension and application of causal isotonic regression to this setting. When it is of interest to estimate the distribution of disease duration under current status sampling, CIR offers a flexible approach accommodating the use of off-the-shelf machine learning tools. Unlike the commonly used nonparametric maximum likelihood estimator, the validity of CIR relies only on conditional independence of survey response times and event indicators within strata of measured baseline covariates, thereby weakening the assumptions under which valid results will be obtained. In addition, we have shown how CIR can be implemented to properly account for survey nonresponse.